# Stable subgroups of graph products

### v1

@@ 2.4.4 Star Cayley graph
The following was observed by Valiunas in [valiunas2021graphproducts, Proof of Corollary D], and follows from [genevois2017, Section 8.1] and [GenevoisMartin, Theorem 2.10]. We include a short proof here for completeness.
@@
[genevois2017, GenevoisMartin] For a graph product $\Gamma\mathcal{G}$, the contact graph of the prism complex is $\Gamma\mathcal{G}$-equivariantly quasi-isometric to the star Cayley graph.
@@ 3 Disk diagrams
Many of the proofs in this paper are visualized using disk diagrams, a version of the dual van Kampen diagrams used for graph products by Valiunas [valiunas2021graphproducts, Section 6.2], and also by Genevois [Genevois_VanKampen]. For finitely presented groups this notion is classical (cf. the pictures of [MR787801]), but our particular use of disk diagrams for graph products closely mirrors their more recent use for right-angled Artin groups ([CrispWiest], [Kim], [KimKoberda2]). In Section 3.1 we review the construction of such diagrams, and in Section 3.2 we note some basic lemmas that will be important to our applications. In Section 3.3, Section 3.4, and Section 3.5 we prove some more technical disk diagram results that will be key to the proof of the main theorem.
@@ 3.1 Constructing disk diagrams
By the nature of the relators for a presentation of $\Gamma\mathcal{G}$ using the prism generating set $S_{p}$, a van Kampen diagram $\mathcal{D}_{\mathsf{vK}}$ for an identity word $s=s_{1}\dots s_{k}$ will be a simply-connected complex made of squares, triangles, and edges embedded in $\mathbb{R}^{2}$, with edges labeled by the letters $s_{i}\in S_{p}$ around the boundary (see, for example, [LS]). Our convention will be to read boundary words clockwise in van Kampen diagrams, their duals, and disk diagrams alike. Observe that (geometric) hyperplanes and their carriers can be defined for $\mathcal{D}_{\mathsf{vK}}$ in the same way as they are for the prism complex, using a cellular map from $\mathcal{D}_{\mathsf{vK}}$ to $P(\Gamma\mathcal{G})$ that respects edge labels. In particular, each geometric hyperplane and hyperplane carrier in a van Kampen diagram is a connected component of the preimage of a geometric hyperplane or hyperplane carrier in the prism complex, and two geometric hyperplanes can only intersect in a van Kampen diagram if their corresponding hyperplanes in the prism complex are transverse.
@@ 3.2 Disk diagram lemmas
Our next lemma, about geodesic subwords of an identity word, recovers as a corollary that two vertices in $\text{Cay}_{p}(\Gamma\mathcal{G})$ have only finitely many geodesics between them (whereas $\text{Cay}_{p}(\Gamma\mathcal{G})$ is locally infinite). This fact is also a direct consequence of the normal form for graph products [greennormalform].
@@
[greennormalform, Theorem 3.9] The $S_{p}$–geodesic word for a given element in $\Gamma\mathcal{G}$ is unique up to ordering of its letters, so there are only finitely many prism geodesics between two vertices in $\text{Cay}_{p}(\Gamma\mathcal{G})$.
@@
[Genevois_VanKampen, Theorem 3.2] Let $w=w_{1}\cdots w_{n}$ be a non-trivial prism word of a graph product $\Gamma\mathcal{G}$. This word is geodesic if and only if for any $i<j$, if $w_{i}$ and $w_{j}$ belong to the same vertex group $G_{v}$, then there exists $i<k<j$ whose associated vertex in $\Gamma$ does not belong to $\mathsf{Star}{(v)}$.
@@ 3.4 Disk diagrams for almost star-free subgroups
The first of this series of lemmas gives a quantitative control on how $S_{H}$–geodesic words reduce to $S_{p}$–geodesic words. It is adapted from Lemma 4.1 in [RAAGstable], a right-angled Artin group counterpart bounding “cancellation diameter”; in RAAG disk diagrams, the analogue for dual graphs are arcs that represent the cancellation of a generator with its inverse, where these are letters from the (finite) standard generating set for the RAAG. In our setting of graph products, dual graphs represent the combining of prism generators from the same vertex group to yield the identity (as observed in 3.8).
@@
Like the previous lemma, the following lemma mirrors a right-angled Artin group version in [RAAGstable], this time of the same name. This concerns vanishing subwords for $S_{H}$–geodesic words $h$, defined as follows. Suppose $h=h_{1}\cdots h_{n}$ where $h_{i}\in S_{H}$ are $S_{p}$–geodesic words and $n$ is minimal. We say the subword $h_{i}\cdots h_{j}$ is vanishing if there exists an $S_{p}$–geodesic word $w$ equal to $h$ with a disk diagram for $h\overline{w}$ where no dual graph rooted in $\overline{w}$ intersects $h_{i}\cdots h_{j}$. In this sense, the subword $h_{i}\cdots h_{j}$ does not contribute to the letters in the prism-geodesic word $w$.
@@ 3.5 Join-busting
This proof works in the same way as [RAAGstable, Lemma 5.1]. ∎
@@
Let $\Gamma$ be a finite simple graph with no isolated vertices. Let $\mathcal{G}=\{G_{v}\,|\,v\in V(\Gamma)\}$ be a collection of finitely generated infinite groups, and $P(\Gamma\mathcal{G})$ the prism complex of the graph product $\Gamma\mathcal{G}$. Let $H$ be a finitely generated subgroup of the graph product $\Gamma\mathcal{G}$. Then the following are equivalent.
@@ 4.1 Almost join-free subgroups embed quasi-isometrically into the contact graph
The proof follows [RAAGstable, Proposition 4.4]. Recall that $|\cdot|_{H}$ is the $S_{H}$–length of an element in $H$, and $|\cdot|_{p}$ the prism-length of an element in $\Gamma\mathcal{G}$. By definition of $\mathsf{m}_{H}$, for every $h\in H$ we have that $|h|_{H}\geq|h|_{p}/\mathsf{m}_{H}$. Now consider a prism-geodesic word $w$ for $h$, and a disk diagram for $h\overline{w}$. By 3.9, any dual graph with a root in $\overline{w}$ has its remaining roots in $h$. By 3.19 each graph dual to a letter of $\overline{w}$ has its extremal roots in $h$ at $S_{H}$–distance at most $D$, and 3.20 ensures there is no vanishing subword of $h$ of $S_{H}$–length greater than $K$. We can see that $|h|_{H}$ is maximal relative to $|h|_{p}=|w|_{p}$ when none of the dual graphs rooted in both $h$ and $w$ cross each other, all of the roots in $h$ of each of these dual graphs are as far apart as possible, and the roots of different dual graphs are separated by maximally large vanishing subwords, as illustrated in Figure 13. This implies that $|h|_{H}\leq(K+D)|h|_{p}+K$. ∎
@@ 4.2 Stable subgroups are almost join-free
Consider a standard generating set $S$ for $\Gamma\mathcal{G}$ (ie. the generating set obtained by taking a union of finite generating sets for each vertex group in $\mathcal{G}$). By the normal form for graph products [greennormalform], under this generating set every parabolic subgroup can be considered as being isometrically embedded in $\Gamma\mathcal{G}$. As $\Lambda$} is a join of two subgraphs $\Gamma_{1}$ and $\Gamma_{2}$ of $\Gamma$, recall that $G_{\Lambda}$ is isomorphic to $G_{\Gamma_{1}}\times G_{\Gamma_{2}}$, and each one of $G_{\Lambda},G_{\Gamma_{1}}$, and $G_{\Gamma_{2}}$ are isometrically embedded in $\Gamma\mathcal{G}$ under the generating set $S$.
As $A=H\cap G_{\Lambda}$ is infinite, the projection of $A$ to at least one of $G_{\Gamma_{1}}$ and $G_{\Gamma_{2}}$ must also be infinite. Call these projections $X$ and $Y$ respectively, and without loss of generality assume that $X$ is infinite. For every $M>0$ we can therefore pick $x_{1},x_{2}\in X$ and $y_{1},y_{2}\in Y$ such that $d_{S}(x_{1},x_{2})>M$ and $x_{1}y_{1},x_{2}y_{2}\in A$. It follows from [DT, Lemma 6.2] that there exist (3,0)-quasi-geodesics in $G_{\Lambda}$, and therefore in $\Gamma\mathcal{G}$, between $x_{1}y_{1}$ and $x_{2}y_{2}$ such that at least one of these is not in the $M$-neighborhood of the other. This means that $H$ cannot be stable. ∎
@@ 4.3 Quasi-isometrically embedded subgroups are stable
If $H$ is finite, then it is automatically stable, so assume that $H$ is infinite. Recall from 2.7 that $\mathcal{C}P(\Gamma\mathcal{G})$ is hyperbolic, and that the action of $\Gamma\mathcal{G}$ on $\mathcal{C}P(\Gamma\mathcal{G})$ is acylindrical. Because the orbit of $H$ into $\mathcal{C}P(\Gamma\mathcal{G})$ is a quasi-isometric embedding by assumption, we can conclude that $(\Gamma\mathcal{G},\mathcal{C}P(\Gamma\mathcal{G}),H)$ is an A/QI triple, as defined in [abbott2022acylindrically, Definition 1.2]. We also have that $\Gamma\mathcal{G}$ is finitely generated by assumption. Thus, [abbott2022acylindrically, Theorem 1.5] implies that $H$ is stable. ∎
@@ 5 Purely loxodromic subgroups
In [RAAGstable, Theorem 1.1], it is shown that the stable subgroups of right-angled Artin groups are exactly the finitely generated purely loxodromic subgroups. We recall that an isometry $g$ of a hyperbolic space $X$ is called loxodromic if $\langle g\rangle$ has unbounded orbits, and fixes exactly two points on the Gromov boundary $\partial X$. If $G$ is a group acting by isometries on $X$, a subgroup $H\leqslant G$ is purely loxodromic if every infinite order $h\in H$ acts loxodromically on $X$.
We can show that every stable subgroup of a graph product is purely loxodromic via the following standard observation, combined with 1.2 and the fact that the action of $\Gamma\mathcal{G}$ on $\mathcal{C}P(\Gamma\mathcal{G})$ is acylindrical (see 2.7). Note that in an acylindrical action on a hyperbolic space, every element with unbounded orbits is loxodromic [Bowditch2008, Lemma 2.2].
@@
However, if we add the assumption that $\Gamma\mathcal{G}$ cannot contain infinite torsion subgroups (for example if $\Gamma\mathcal{G}$ is virtually torsion-free, as in the right-angled Artin group case), we are able to obtain equivalence. The following proposition (which will help us establish the equivalence) is known, see for example [Fioravanti2024, Proposition 2.10], however we include a short proof here for completeness.
@@
As a result, in this case we obtain a fourth condition equivalent to stability.
@@
Let $\Gamma$ be a finite simple graph with no isolated vertices, $\mathcal{G}=\{G_{v}\,|\,v\in V(\Gamma)\}$ be a collection of infinite finitely generated groups with no infinite torsion subgroups, and $P(\Gamma\mathcal{G})$ the prism complex of the graph product $\Gamma\mathcal{G}$. Let $H$ be a finitely generated subgroup of the graph product $\Gamma\mathcal{G}$. Then the following are equivalent.
  1. 1.
    $H$ is stable in $\Gamma\mathcal{G}$.
  2. 2.
  3. 3.
    $H$ is almost join-free.
  4. 4.
    $H$ is purely loxodromic with respect to the action on $\mathcal{C}P(\Gamma\mathcal{G})$.

### v2

@@ 2.4.4 Star Cayley graph
The following was observed by Valiunas in [38, Proof of Corollary D], and follows from [17, Section 8.1] and [16, Theorem 2.10]. We include a short proof here for completeness.
@@
[17, 16] For a graph product $\Gamma\mathcal{G}$, the contact graph of the prism complex is $\Gamma\mathcal{G}$-equivariantly quasi-isometric to the star Cayley graph.
@@ 3 Disk diagrams
Many of the proofs in this paper are visualized using disk diagrams, a version of the dual van Kampen diagrams used for graph products by Valiunas [38, Section 6.2], and also by Genevois [19]. For finitely presented groups this notion is classical (cf. the pictures of [14]), but our particular use of disk diagrams for graph products closely mirrors their more recent use for right-angled Artin groups ([11], [27], [26]). In Section 3.1 we review the construction of such diagrams, and in Section 3.2 we note some basic lemmas that will be important to our applications. In Section 3.3, Section 3.4, and Section 3.5 we prove some more technical disk diagram results that will be key to the proof of the main theorem.
@@ 3.1 Constructing disk diagrams
By the nature of the relators for a presentation of $\Gamma\mathcal{G}$ using the prism generating set $S_{p}$, a van Kampen diagram $\mathcal{D}_{\mathsf{vK}}$ for an identity word $s=s_{1}\dots s_{k}$ will be a simply-connected complex made of squares, triangles, and edges embedded in $\mathbb{R}^{2}$, with edges labeled by the letters $s_{i}\in S_{p}$ around the boundary (see, for example, [29]). Our convention will be to read boundary words clockwise in van Kampen diagrams, their duals, and disk diagrams alike. Observe that (geometric) hyperplanes and their carriers can be defined for $\mathcal{D}_{\mathsf{vK}}$ in the same way as they are for the prism complex, using a cellular map from $\mathcal{D}_{\mathsf{vK}}$ to $P(\Gamma\mathcal{G})$ that respects edge labels. In particular, each geometric hyperplane and hyperplane carrier in a van Kampen diagram is a connected component of the preimage of a geometric hyperplane or hyperplane carrier in the prism complex, and two geometric hyperplanes can only intersect in a van Kampen diagram if their corresponding hyperplanes in the prism complex are transverse.
@@ 3.2 Disk diagram lemmas
Our next lemma, about geodesic subwords of an identity word, recovers as a corollary that two vertices in $\text{Cay}_{p}(\Gamma\mathcal{G})$ have only finitely many geodesics between them (whereas $\text{Cay}_{p}(\Gamma\mathcal{G})$ is locally infinite). This fact is also a direct consequence of the normal form for graph products [20].
@@
[20, Theorem 3.9] The $S_{p}$–geodesic word for a given element in $\Gamma\mathcal{G}$ is unique up to ordering of its letters, so there are only finitely many prism geodesics between two vertices in $\text{Cay}_{p}(\Gamma\mathcal{G})$.
@@
[19, Theorem 3.2] Let $w=w_{1}\cdots w_{n}$ be a non-trivial prism word of a graph product $\Gamma\mathcal{G}$. This word is geodesic if and only if for any $i<j$, if $w_{i}$ and $w_{j}$ belong to the same vertex group $G_{v}$, then there exists $i<k<j$ whose associated vertex in $\Gamma$ does not belong to $\mathsf{Star}{(v)}$.
@@ 3.4 Disk diagrams for almost star-free subgroups
The first of this series of lemmas gives a quantitative control on how $S_{H}$–geodesic words reduce to $S_{p}$–geodesic words. It is adapted from Lemma 4.1 in [28], a right-angled Artin group counterpart bounding “cancellation diameter”; in RAAG disk diagrams, the analogue for dual graphs are arcs that represent the cancellation of a generator with its inverse, where these are letters from the (finite) standard generating set for the RAAG. In our setting of graph products, dual graphs represent the combining of prism generators from the same vertex group to yield the identity (as observed in 3.8).
@@
Like the previous lemma, the following lemma mirrors a right-angled Artin group version in [28], this time of the same name. This concerns vanishing subwords for $S_{H}$–geodesic words $h$, defined as follows. Suppose $h=h_{1}\cdots h_{n}$ where $h_{i}\in S_{H}$ are $S_{p}$–geodesic words and $n$ is minimal. We say the subword $h_{i}\cdots h_{j}$ is vanishing if there exists an $S_{p}$–geodesic word $w$ equal to $h$ with a disk diagram for $h\overline{w}$ where no dual graph rooted in $\overline{w}$ intersects $h_{i}\cdots h_{j}$. In this sense, the subword $h_{i}\cdots h_{j}$ does not contribute to the letters in the prism-geodesic word $w$.
@@ 3.5 Join-busting
This proof works in the same way as [28, Lemma 5.1]. ∎
@@
Let $\Gamma$ be a finite simple graph with no isolated vertices, $\mathcal{G}=\{G_{v}\,|\,v\in V(\Gamma)\}$ be a collection of finitely generated infinite groups, and $P(\Gamma\mathcal{G})$ be the prism complex of the graph product $\Gamma\mathcal{G}$. Let $H$ be a finitely generated subgroup of the graph product $\Gamma\mathcal{G}$. Then the following are equivalent.
@@ 4.1 Almost join-free subgroups embed quasi-isometrically into the contact graph
The proof follows [28, Proposition 4.4]. Recall that $|\cdot|_{H}$ is the $S_{H}$–length of an element in $H$, and $|\cdot|_{p}$ the prism-length of an element in $\Gamma\mathcal{G}$. By definition of $\mathsf{m}_{H}$, for every $h\in H$ we have that $|h|_{H}\geq|h|_{p}/\mathsf{m}_{H}$. Now consider a prism-geodesic word $w$ for $h$, and a disk diagram for $h\overline{w}$. By 3.9, any dual graph with a root in $\overline{w}$ has its remaining roots in $h$. By 3.19 each graph dual to a letter of $\overline{w}$ has its extremal roots in $h$ at $S_{H}$–distance at most $D$, and 3.20 ensures there is no vanishing subword of $h$ of $S_{H}$–length greater than $K$. We can see that $|h|_{H}$ is maximal relative to $|h|_{p}=|w|_{p}$ when none of the dual graphs rooted in both $h$ and $w$ cross each other, all of the roots in $h$ of each of these dual graphs are as far apart as possible, and the roots of different dual graphs are separated by maximally large vanishing subwords, as illustrated in Figure 13. This implies that $|h|_{H}\leq(K+D)|h|_{p}+K$. ∎
@@ 4.2 Stable subgroups are almost join-free
Consider a standard generating set $S$ for $\Gamma\mathcal{G}$ (i.e. the generating set obtained by taking a union of finite generating sets for each vertex group in $\mathcal{G}$). By the normal form for graph products [20], under this generating set every parabolic subgroup can be considered as being isometrically embedded in $\Gamma\mathcal{G}$. As $\Lambda$} is a join of two subgraphs $\Gamma_{1}$ and $\Gamma_{2}$ of $\Gamma$, recall that $G_{\Lambda}$ is isomorphic to $G_{\Gamma_{1}}\times G_{\Gamma_{2}}$, and each one of $G_{\Lambda},G_{\Gamma_{1}}$, and $G_{\Gamma_{2}}$ are isometrically embedded in $\Gamma\mathcal{G}$ under the generating set $S$.
As $A=H\cap G_{\Lambda}$ is infinite, the projection of $A$ to at least one of $G_{\Gamma_{1}}$ and $G_{\Gamma_{2}}$ must also be infinite. Call these projections $X$ and $Y$ respectively, and without loss of generality assume that $X$ is infinite. For every $M>0$ we can therefore pick $x_{1},x_{2}\in X$ and $y_{1},y_{2}\in Y$ such that $d_{S}(x_{1},x_{2})>M$ and $x_{1}y_{1},x_{2}y_{2}\in A$. It follows from [12, Lemma 6.2] that there exist (3,0)-quasi-geodesics in $G_{\Lambda}$, and therefore in $\Gamma\mathcal{G}$, between $x_{1}y_{1}$ and $x_{2}y_{2}$ such that at least one of these is not in the $M$-neighborhood of the other. This means that $H$ cannot be stable. ∎
@@ 4.3 Quasi-isometrically embedded subgroups are stable
If $H$ is finite, then it is automatically stable, so assume that $H$ is infinite. Recall from 2.7 that $\mathcal{C}P(\Gamma\mathcal{G})$ is hyperbolic, and that the action of $\Gamma\mathcal{G}$ on $\mathcal{C}P(\Gamma\mathcal{G})$ is acylindrical. Because the orbit of $H$ into $\mathcal{C}P(\Gamma\mathcal{G})$ is a quasi-isometric embedding by assumption, we can conclude that $(\Gamma\mathcal{G},\mathcal{C}P(\Gamma\mathcal{G}),H)$ is an A/QI triple, as defined in [2, Definition 1.2]. We also have that $\Gamma\mathcal{G}$ is finitely generated by assumption. Thus, [2, Theorem 1.5] implies that $H$ is stable. ∎
@@ 5 Purely loxodromic subgroups
In [28, Theorem 1.1], it is shown that the stable subgroups of right-angled Artin groups are exactly the finitely generated purely loxodromic subgroups. We recall that an isometry $g$ of a hyperbolic space $X$ is called loxodromic if $\langle g\rangle$ has unbounded orbits, and fixes exactly two points on the Gromov boundary $\partial X$. If $G$ is a group acting by isometries on $X$, a subgroup $H\leqslant G$ is purely loxodromic if every infinite order $h\in H$ acts loxodromically on $X$.
We can show that every stable subgroup of a graph product is purely loxodromic via the following standard observation, combined with 1.2 and the fact that the action of $\Gamma\mathcal{G}$ on $\mathcal{C}P(\Gamma\mathcal{G})$ is acylindrical (see 2.7). Note that in an acylindrical action on a hyperbolic space, every element with unbounded orbits is loxodromic [7, Lemma 2.2].
@@
However, if we add the assumption that $\Gamma\mathcal{G}$ cannot contain infinite torsion subgroups (for example if $\Gamma\mathcal{G}$ is virtually torsion-free, as in the right-angled Artin group case), we are able to obtain equivalence. The following proposition (which will help us establish the equivalence) is known, see for example [15, Proposition 2.10], however we include a short proof here for completeness.
@@
As a result, by combining 1.2 with 5.1 and 5.3, in this case we obtain a additional condition equivalent to stability.
Corollary 5.4.
Let $\Gamma$ be a finite simple graph with no isolated vertices, $\mathcal{G}=\{G_{v}\,|\,v\in V(\Gamma)\}$ be a collection of infinite finitely generated groups with no infinite torsion subgroups, and $P(\Gamma\mathcal{G})$ the prism complex of the graph product $\Gamma\mathcal{G}$. A finitely generated subgroup $H$ of $\Gamma\mathcal{G}$ is stable if and only if it is purely loxodromic with respect to the action on the contact graph $\mathcal{C}P(\Gamma\mathcal{G})$.
6 Infinite index Morse subgroups
A subgroup $H$ of a finitely generated group $G$ is said to be Morse (sometimes called strongly quasiconvex) if for any constants $\lambda\geqslant 1,c\geqslant 0$, there is a constant $M\geqslant 0$ (depending on $\lambda,c$) such that every $(\lambda,c)$–quasi-geodesic in $G$ with endpoints in $H$ is contained in the MM–neighborhood of $H$. Morse subgroups can be related to stable subgroups via the following result.
Proposition 6.1.
[36, Theorem 4.8][18] Let $G$ be a finitely generated group. A subgroup $H$ of $G$ is stable if and only if it is Morse and hyperbolic.
In [36], it is shown that the stable subgroups of right-angled Artin groups with connected defining graphs are exactly the infinite index Morse subgroups. This was also shown for mapping class groups in [24]. Here we extend this characterization to the stable subgroups of graph products of infinite groups, where the defining graph is connected. This connected assumption is necessary, as the factors in a free product will always be Morse and infinite index, but not necessarily hyperbolic.
We begin with the following easy observation.
Lemma 6.2.
Let $\Gamma\mathcal{G}$ be a graph product of finitely generated infinite groups, where $\Gamma$ has no isolated vertices. If $H\leqslant\Gamma\mathcal{G}$ is stable, then $H$ is infinite index and Morse.
Proof.
It follows from 6.1 that $H$ is Morse. As $\Gamma$ has no isolated vertices, $\Gamma\mathcal{G}$ contains a (quasi-isometrically embedded) direct product of infinite groups, so is not hyperbolic, so no finite index subgroup is hyperbolic. Again using 6.1 (or [12, Lemma 3.3]), this means that no finite index subgroup of $\Gamma\mathcal{G}$ is stable, which gives us the conclusion. ∎
The proof of the reverse direction of 6.2 (in the case that $\Gamma$ is connected) uses the fact that Morse subgroups have finite height. A subgroup $H$ of a group $G$ is said to have finite height in $G$ if there exists $k$ such that, for any $g_{1},\dots,g_{k}\in G$ where the cosets $g_{1}H,\ldots,g_{k}H$ are distinct, the intersection of the conjugates $g_{1}Hg_{1}^{-1},\ldots,g_{k}Hg_{k}^{-1}$ is finite.
@@ 6 Infinite index Morse subgroups
[36, Theorem 4.15] If $H$ is a Morse subgroup of a finitely generated group $G$, then $H$ has finite height in $G$.
The following results are adapted from the statements for right-angled Artin groups in [36]. The next proposition is well known to experts, the proof is included for completeness.
Proposition 6.4.
If $G_{1}$ and $G_{2}$ are finitely generated infinite groups, and $H\leqslant G_{1}\times G_{2}$ is a Morse subgroup, then either $H$ is finite, or $H$ has finite index in $G_{1}\times G_{2}$.
Proof.
There are two cases to consider. The first case is that at least one of $H\cap G_{1}$ or $H\cap G_{2}$ is infinite. Suppose without loss of generality that $H\cap G_{2}$ is infinite. We want to show that $H$ has finite index in $G_{1}\times G_{2}$. For this purpose, it is enough to show that both $G_{1}:H\cap G_{1}]$ and $G_{2}:H\cap G_{2}]$ are finite.
Suppose for the purpose of contradiction that $H\cap G_{1}$ does not have finite index in $G_{1}$. This means that for any $k\geqslant 1$, we can find $a_{1},\ldots,a_{k}\in G_{1}$ such that $a_{1}(H\cap G_{1}),\ldots,a_{k}(H\cap G_{1})$ are distinct. This implies that $a_{1}H,\ldots,a_{k}H$ are also distinct. Each $a_{i}$ commutes with the elements of $G_{2}$, so $H\cap G_{2}\leqslant a_{i}Ha_{i}^{-1}$, and therefore $H\cap G_{2}\leq\bigcap_{i=1}^{k}a_{i}Ha_{i}^{-1}$. As $H\cap G_{2}$ is infinite, $H$ does not have finite height, contradicting 6.3.
We therefore have that $H\cap G_{1}$ has finite index in $G_{1}$. As $G_{1}$ is infinite, this implies that $H\cap G_{1}$ is also infinite, so by the same reasoning as in the previous paragraph $H\cap G_{2}$ also has finite index in $G_{2}$. We can therefore conclude that $H$ has finite index in $G_{1}\times G_{2}$.
The second case is that both $H\cap G_{1}$ and $H\cap G_{2}$ are finite. In this case we will show that $H$ is finite, so suppose for the purpose of contradiction that $H$ is infinite. We must therefore have that the projection of $H$ to at least one of $G_{1}$ or $G_{2}$ is infinite. Suppose without loss of generality that the projection to $G_{1}$ is infinite.
Let $S_{1}$ and $S_{2}$ be finite generating sets for $G_{1}$ and $G_{2}$ respectively. As $H\cap G_{1}$ is finite, for any $b\in G_{2}$ we have that the set $G_{1}\times\{b\}$ contains only finitely many elements of $H$. It follows that, for any $M\geqslant 1$, the set $G_{1}\times B_{S_{2}}(M)$ contains only finitely many elements of $H$, where $B_{S_{2}}(M)$ is the ball of radius $M$ around $\mathsf{id}_{G_{2}}$ in $G_{2}$ with respect to $S_{2}$. Now let $h_{1}h_{2}\in H$ such that $h_{1}\in G_{1}$ and $h_{2}\in G_{2}$, and let $\gamma_{1}$ be an $S_{1}$–geodesic from $\mathsf{id}_{G_{1}}$ to $h_{1}$, and $\gamma_{2}$ be an $S_{2}$–geodesic from $\mathsf{id}_{G_{2}}$ to $h_{2}$. The concatenation of $\gamma_{1}\times\{\mathsf{id}_{G_{2}}\}$ with $\{h_{1}\}\times\gamma_{2}$ is therefore a geodesic in $G_{1}\times G_{2}$ from $\mathsf{id}_{G_{1}\times G_{2}}$ to $h_{1}h_{2}$, with both of these points lying in $H$.
As the projection of $H$ to $G_{1}$ is infinite, we can choose $h_{1}$ to have arbitrarily large $S_{1}$–length. As $G_{1}\times B_{S_{2}}(M)$ contains only finitely many elements of $H$, say $m$ many, we choose $h_{1}$ such that $|h_{1}|_{S_{1}}\geqslant 3M(m+1)$. Therefore some point in $\gamma_{1}\times\{\mathsf{id}_{G_{2}}\}$ has distance at least $M$ from any point of $H$, which contradicts $H$ being Morse. Therefore $H$ must be finite. ∎
Lemma 6.5.
Let $\Gamma\mathcal{G}$ be a graph product of infinite groups, with $\Gamma$ connected. For any $v\in V(\Gamma)$, and any $g_{1},g_{2}\in\Gamma\mathcal{G}$, there is a finite sequence of conjugates of star subgroups $g_{1}G_{\mathsf{Star}(v)}g_{1}^{-1}=Q_{0},Q_{1},\ldots,Q_{m}=g_{2}G_{\mathsf{Star}(v)}g_{2}^{-1}$ such that $Q_{i-1}\cap Q_{i}$ is infinite for each $i\in\{1,\ldots,m\}$.
Proof.
This follows the proof of Lemma 8.17 in [36], with the induction done on $n=|g_{1}^{-1}g_{2}|_{\star}$. ∎
Corollary 6.6.
Let $\Gamma\mathcal{G}$ be a graph product of infinite groups, with $\Gamma$ connected. For any join subgraph $\Lambda$ of $\Gamma$, and any $g_{1},g_{2}\in\Gamma\mathcal{G}$, there is a finite sequence of conjugates of join subgroups $g_{1}G_{\Lambda}g_{1}^{-1}=Q_{0},Q_{1},\ldots,Q_{m}=g_{2}G_{\Lambda}g_{2}^{-1}$ such that $Q_{i-1}\cap Q_{i}$ is infinite for each $i\in\{1,\ldots,m\}$.
Proof.
Take any $v\in V(\Lambda)$, then $G_{\mathsf{Star}(v)}\cap G_{\Lambda}$ is infinite. We can therefore take $Q_{1}=g_{1}G_{\mathsf{Star}(v)}g_{1}^{-1}$ and $Q_{m-1}=g_{2}G_{\mathsf{Star}(v)}g_{2}^{-1}$, and as star subgroups are join subgroups, the conclusion follows from 6.5. ∎
Corollary 6.7.
Let $\Gamma\mathcal{G}$ be a graph product of finitely generated infinite groups, with $\Gamma$ connected. If $H\leqslant\Gamma\mathcal{G}$ is infinite index and Morse, then $H$ is almost join-free.
Proof.
This follows the proof of Proposition 8.18 in [36], using 6.4 in place of [36, Lemma 8.16], 6.6 in place of [36, Lemma 8.17], join subgroups in place of star subgroups, finite intersections in place of trivial ones, and infinite intersections in place of non-trivial ones. ∎
We finally combine 1.2, 6.2, and 6.7 to obtain the following.
Corollary 6.8.
Let $\Gamma\mathcal{G}$ be a graph product of finitely generated infinite groups, with $\Gamma$ a connected graph with at least two vertices. A subgroup $H$ of $\Gamma\mathcal{G}$ is stable if and only if it is infinite index and Morse.